What is the resistance of a silver wire 25 m long at 20 °C and whose diameter is 0.00095 cm? the of silver is at 20 °C is 1.6 x 10-8Ω.m.


1
Expert's answer
2021-11-25T10:42:01-0500

We solve the problem using the formula

"R=\frac{\rho\times L}{A}"

We have the parameters:

Length of wire, "l" ="25m"

diameter of wire, "d=0.00095cm"

Resistivity, "\rho=1.6\times10^{-8} Ωm"


We convert diameter from "cm" to "m" and then divide the results by 2 to get our "radius"

"d=\frac{0.00095}{100}m=0.0000095m"

Radius, "r=\frac{0.0000095}{2}=4.75\times10^{-6}m"


Therefore, we get the value for "Area" by the formula

"A=\pi\times r²"

"A=\frac{22}{7}\times[4.75\times10^-{6}]²"

"A=\frac{22}{7}\times2.256\times10^{-11}"

"A=7.09\times10^{-11}m²"


Now, we have the necessary parameters to work with which are, "Length (m)" ,"Resistivity,(\rho)" and "Area(m\u00b2)"

We then get the resistance of the wire using the formula

"R=\frac{\rho\times L}{A}"

"R=\frac{1.6\times10^{-8}\times25}{7.09\times10^{-11}}"


"R=\frac{4\times10^{-7}}{7.09\times10^{-11}}"

"R=5.64\times10^{-19}Ω"

Therefore, the resistance of the silver wire is

"5.64\times10^{-19}Ω"